What power lens is needed to correct for the farsightedness of a person whose near point is 58.0 cm? (Assume that eye-lens distance is negligible and the normal near point of the eye is 25.0 cm). Answer in diopters please.
Answer #1

Assuming the eye lens distance to be negligible. We want that when an object is at 25cm from the person’s eyes, the spectacle lens must produce an image 58.0 cm away (the near point).

So the image distance is di = -58.0 cm = -0.58 m

The image distance is negative, because it is on the same side of the spectacle as the object.

And the object distance is do = 25cm = 0.25m

Power of lens required by the person

P=\frac{1}{d_{o}}+\frac{1}{d_{i}} = \frac{1}{0.25}+\frac{1}{-0.58} = +2.276

Positive powerof lens means it is of converging type.